How can we use the net present value rule in the following cases a Choose between 2 project with different lives b)ln case of buying new equipment cDelay or postpone the inv estment
Answer #1

Ans Use the Net present value rule

Net present value: it means present value of future cash flow minus present value of cash outflow

a) in case of choose between 2 project with different lives, first we to calculate net present value normally for the project than after that we will calculate equivalent cash flow per year.

Example : if project A NPV is $100 and 5yrs project life

Project B NPV is $150 for 8yrs project life

Discount rate 10% , cumulative value for 5yrs = 3.790

cumulative value for 8yrs = 4.8684

Equivalent cash flow for project A = 100/3.790 = $26.38

Equivalent cash flow for project B = 120/4.8684 = $24.65

in this case we should choose Project A because it is providing more cash flow as compare of project B

b) in Cash buying new equipment: in this case we have calculate present value of future cash and present value of cash outflow if difference is positive is project is acceptable otherwise not

means present value of future cash which is generated by new equipment is more than cash outflow than the project is acceptable.

NPV = PV of cash inflow -PV of cashoutflow

c) NPV : for delay or postpone project : in this case we should check what is the npv of delay project and what is npv of without delay or postpone project . some time npv of delay project is more worth as compare of npv of without delay project. NPV for delay period is always Zero

If present value of cash inflow minus present value of cash outflow is positive than project is acceptable